40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 111     Entry time: Tue Apr 16 00:40:45 2013
Author: Koji 
Type: Optics 
Category: General 
Subject: PD/QPD path gluing ~ preparation 

[Jeff Koji]

- Placed the optics on the PD/QPD path

- Checked the alignment of the beam on the dummy PD/QPD mounts

- There is a bit of (~0.5mm) shift of the spot position from the center. Mainly downward. This is well within a ball park of the PD mounts.

- The PD/QPD path gluing will take place tomorrow.


- Went to the 40m and received the DCPDs from Bob's lab.

- Took six ISC QPDs for the sake of the OMCs.

- They are now in the OMC lab. 


- Measured the B mirror / E mirror R&Ts.

- Found anomalously high loss (3%) for the B mirrors (BSs)

- Went through the all mirrors. Some mirrors (3 or 4) seemed less lossy (<~1%). They will be used for the DCPD BS.

ELOG V3.1.3-